# Magnetic Structures Database from Symmetry-aided High-Throughput Calculations

### v2

@@ III Magnetic Topological Materials
For decades, the topology of electronic bands in condensed-matter materials has been a research hotspot Hasan and Kane (2010), leading to a systematic classification of several tens of thousands non-magnetic topological materials based on the symmetry analysis Zhang et al. (2019); Vergniory et al. (2019); Tang et al. (2019). Owing to the limited number of known magnetic structures Gallego et al. (2016), theoretical predictions of magnetic topological materials remain relatively scarce, with only a few hundred reported to date Xu et al. (2020); Bernevig et al. (2022). Based on our magnetic structure database, we identify 1070 compounds as magnetic topological materials using magnetic topological quantum chemistry Kruthoff et al. (2017); Elcoro et al. (2021). They can be broadly classified into 125 magnetic topological insulators and 945 magnetic topological semimetals, as summarized in Fig. 1b. A complete list of materials and symmetry indicators is provided in the Supplementary Materials 2. To illustrate the physical richness enabled by the symmetry-resolved magnetic structures, we select several representative materials shown below. We also tested the effects of varying the Hubbard $U$ within reasonable ranges, and found that both their magnetic structures and topological properties remain unchanged.
@@ III.1 Magnetic topological insulators
AFM axion insulator FeSe: The FeSe compound (icsd_166445) studied here crystallizes in the orthorhombic space group $Pnma$ (No. 62) Margadonna et al. (2009), which is different from the widely studied superconducting $\alpha$-FeSe of the tetragonal structure $P4/nmm$ (No. 129) Zhang et al. (2018); Wang et al. (2018). Our calculations predict an antiferromagnetic ground state in FeSe. It belongs to the magnetic space group $Pnma$ (62.441), with magnetic moments oriented along the [010] direction, as shown in Fig. 2a. Due to the presence of inversion symmetry, symmetry indicators ($\eta_{4I}$, $z_{2I,1}$, $z_{2I,2}$, $z_{2I,3}$) can be defined using the parities at high-symmetry points Elcoro et al. (2021). $\eta_{4I}$ is defined as the total number of odd-parity states at the eight time-reversal-invariant momenta (TRIMs) modulo 4, while $z_{2I,i}$ is defined as the total number of odd-parity states at the four TRIMs on the $k_{i}=\pi$ plane modulo 2. Our calculated symmetry indicators ($\eta_{4I}$, $z_{2I,1}$, $z_{2I,2}$, $z_{2I,3}$) are (2000). Furthermore, we find that all plane Chern numbers on the TRIM planes vanish for FeSe, and the system exhibits a direct band gap no smaller than 20 meV throughout the entire Brillouin zone, see Fig. 2b. The nontrivial symmetry indicators, together with a direct band gap throughout the entire Brillouin zone, indicate that FeSe is an axion insulator, in which a topological magnetoelectric effect is expected to occur Wan et al. (2011, 2012).
@@ IV Altermagnets
Altermagnets represent a recently identified magnetic class distinct from conventional antiferromagnets, garnering significant attention due to their unique electronic band structures and chiral magnon splitting  Šmejkal et al. (2022b, a); Xu et al. (2025).
In our database, we have identified 392 altermagnetic materials, many of which exhibit pronounced electronic spin splitting and chiral magnon splitting, with a complete list provided in the Supplementary Materials 2. Here we highlight a representative example, CsV2S2O.
CsV2S2O crystallizes in space group No. 123, and its magnetic moments are aligned along [001] direction, see Fig. 4(a). It exhibits strong altermagnetic spin splitting along high-symmetry paths [Fig. 4(b)], reaching a maximum of approximately 1.7 eV near the X point, indicative of a pronounced electronic altermagnetic effect. The magnon spectrum shows a chiral splitting of up to 94.4 meV at the X point in the Brillouin zone, originating from two symmetry-inequivalent next-nearest-neighbor exchange interactions 2, in contrast to the typical origin which arises from disparities in long-range exchange interactions (e.g., $J_{7a}$ versus $J_{7b}$ in MnF2 Morano et al. (2025)). This value significantly exceeds the experimentally reported chiral magnon splittings in MnF2 Morano et al. (2025) and MnTe Liu et al. (2024).
@@ V Outlook
We have developed a highly efficient methodology that addresses a bottleneck in magnetic materials research by combinating symmetry analysis with first-principles calculations. For continuous magnetic phase transitions, the symmetry of the ordered phase follows a subgroup relation with respect to the parent space group. Using the Wyckoff positions of magnetic atoms, we construct the magnetic representation and decompose it into the irreducible representations of the parent crystallographic space group, thereby reducing the space of candidate MSGs to those compatible with the corresponding irreducible representations in Eq. (1). As the irreps and the corresponding MSGs do not uniquely fix the magnetic structure, we employ the symmetry-adapted basis vectors of the irreducible representation, which provide a more proper strategy for constructing explicit magnetic configurations, yet residual degrees of freedom remain. We further resolve these ambiguities by applying the maximal-symmetry criterion, and by aligning magnetic moments along all the highest-symmetry crystallographic directions. In this way, our approach yields a group theory based framework that drastically reduces the initial input magnetic configuration space and thereby enables efficient high-throughput first-principles calculations.
To empirically assess the validity of our developed symmetry-adapted approach, we analyzed the MAGNDATA database for materials with $\mathbf{q}=0$ and fewer than 40 atoms per primitive cell. Among the 328 materials examined, only 26 out of 328 are incompatible with Landau theory. For the remaining materials, deviations of magnetic moments from high-symmetry directions are found in only 23 out of 302 cases. These results show that the assumptions underlying our symmetry-based framework of using maximal symmetry criterion and aligning basis vector of irrep along all the highest-symmetry directions apply to the vast majority of known magnetic materials (i.e. 92%). When combined with first-principles calculations, the benchmark achieves an accuracy of about 70% (212 out of 302), with the remaining discrepancies mainly due to limitations of first-principles energetics. The applicability of our method is well justified by benchmarking tests.
@@
The efficiency of our framework has enabled the construction of a comprehensive database of magnetic structures with $\mathbf{q}=0$ and fewer than 20 atoms per primitive cell. Both constraints can be straightforwardly relaxed to incorporate commensurate propagation vectors and more atoms in unit cell. Extending the approach to two-dimensional magnetic materials is both natural and highly relevant, given the rapidly growing interest in low-dimensional magnetism. Beyond the topological classification already accomplished, we are actively computing additional physical properties—including multiferroic responses, anomalous Hall conductivity, transport coefficients, etc.—for the magnetic materials identified in this work. These extensions are currently underway.

### v3

@@ III Magnetic Topological Materials
For decades, the topology of electronic bands in condensed-matter materials has been a research hotspot Hasan and Kane (2010), leading to a systematic classification of several tens of thousands non-magnetic topological materials based on the symmetry analysis Zhang et al. (2019); Vergniory et al. (2019); Tang et al. (2019). Owing to the limited number of known magnetic structures Gallego et al. (2016), theoretical predictions of magnetic topological materials remain relatively scarce, with only a few hundred reported to date Xu et al. (2020); Bernevig et al. (2022). Based on our magnetic structure database, we identify 1070 compounds as magnetic topological materials using magnetic topological quantum chemistry Kruthoff et al. (2017); Elcoro et al. (2021). They can be broadly classified into 125 magnetic topological insulators and 945 magnetic topological semimetals, as summarized in Fig. 1b. A complete list of materials and symmetry indicators is provided in the Supplementary Materials SM . To illustrate the physical richness enabled by the symmetry-resolved magnetic structures, we select several representative materials shown below. We also tested the effects of varying the Hubbard $U$ within reasonable ranges, and found that both their magnetic structures and topological properties remain unchanged.
@@ III.1 Magnetic topological insulators
AFM axion insulator FeSe: The FeSe compound (icsd_166445) studied here crystallizes in the orthorhombic space group $Pnma$ (No. 62) Margadonna et al. (2009), which is different from the widely studied superconducting $\alpha$-FeSe of the tetragonal structure $P4/nmm$ (No. 129) Zhang et al. (2018); Wang et al. (2018). Our calculations predict an antiferromagnetic ground state in FeSe. It belongs to the magnetic space group $Pnma$ (62.441), with magnetic moments oriented along the [010] direction, as shown in Fig. 2a. Due to the presence of inversion symmetry, symmetry indicators ($\eta_{4I}$, $z_{2I,1}$, $z_{2I,2}$, $z_{2I,3}$) can be defined using the parities at high-symmetry points Elcoro et al. (2021). $\eta_{4I}$ is defined as the total number of odd-parity states at the eight time-reversal-invariant momenta (TRIMs) modulo 4, while $z_{2I,i}$ is defined as the total number of odd-parity states at the four TRIMs on the $k_{i}=\pi$ plane modulo 2. Our calculated symmetry indicators ($\eta_{4I}$, $z_{2I,1}$, $z_{2I,2}$, $z_{2I,3}$) are (2000). Furthermore, we find that all plane Chern numbers on the TRIM planes vanish for FeSe, and the system exhibits a direct band gap larger than 20 meV throughout the entire Brillouin zone, as shown in Fig. 2b. The nontrivial symmetry indicators, together with a direct band gap throughout the entire Brillouin zone, indicate that FeSe is an axion insulator, in which a topological magnetoelectric effect is expected to occur Essin et al. (2009); Wan et al. (2012).
@@ IV Altermagnets
Altermagnets represent a recently identified magnetic class distinct from conventional antiferromagnets, garnering significant attention due to their unique electronic band structures, chiral magnon splitting, etc. Šmejkal et al. (2022a, b); Xu et al. (2025).
In our database, we have identified 392 altermagnetic materials, many of which exhibit large electronic spin splitting and chiral magnon splitting, with a complete list provided in the Supplementary Materials SM . Here we highlight a representative example, CsV2S2O.
CsV2S2O (icsd_430299) crystallizes in space group $P4/mmm$ (No. 123), and its magnetic moments are aligned along [001] direction, see Fig. 4(a). It exhibits strong altermagnetic spin splitting along high-symmetry paths [Fig. 4(b)], reaching a maximum of approximately 1.7 eV near the X point, indicative of a pronounced electronic altermagnetic effect. The magnon spectrum shows a chiral splitting of up to 94.4 meV at the X point in the Brillouin zone, originating from two symmetry-inequivalent next-nearest-neighbor exchange interactions SM , in contrast to the typical origin which arises from disparities in long-range exchange interactions (e.g., $J_{7a}$ versus $J_{7b}$ in MnF2 Morano et al. (2025)). This value significantly exceeds the experimental chiral magnon splittings in MnF2 Morano et al. (2025) and MnTe Liu et al. (2024).
@@ V Outlook
We have developed a highly efficient methodology that addresses a bottleneck in magnetic materials research by combinating symmetry analysis with first-principles calculations. For continuous magnetic phase transitions, the symmetry of the ordered phase follows a subgroup relation with respect to the parent space group. Using the Wyckoff positions of magnetic atoms, we construct the magnetic representation and decompose it into the irreducible representations of the parent crystallographic space group, thereby reducing the space of candidate MSGs to those compatible with the corresponding irreducible representations in Eq. (1). As the irreps and the corresponding MSGs do not uniquely fix the magnetic structure, we employ the symmetry-adapted basis vectors of the irreducible representation, which provide a more proper strategy for constructing explicit magnetic configurations, yet residual degrees of freedom remain. We further resolve these ambiguities by applying the maximal-symmetry criterion, and by keeping all the basis vectors of irreducible representations compatible with all possible highest-symmetry crystallographic directions. In this way, our strategy yields a group theory based framework that drastically reduces the initial magnetic configuration space, usually $\sim 20$ candidates, and thereby enables efficient high-throughput first-principles calculations.
To assess the scope of our irreducible representation basis vector based strategy, we analyzed 302 materials in the MAGNDATA database. For all materials whose magnetic structures arise from multidimensional irreducible representations ($j>1$ in Eq. (1)), the experimental magnetic structures are fully captured by the maximal-symmetry criterion. By construction, our strategy reduces the otherwise large space of magnetic configurations to compact candidate sets, typically consisting of only tens of predefined configurations per material. Remarkably, these minimal input sets already contain the experimental magnetic structures for 253 compounds (83.8%; see Supplementary Materials SM ). The remaining cases fall outside the scope of our framework, as their magnetic moments deviate from high-symmetry crystallographic directions. Starting from this strongly reduced configuration space, first-principles calculations correctly identify the magnetic structure in 198 of the 253 materials, corresponding to an accuracy of 78.2%, which is expected to further improve with advances in electronic-structure methods.
@@
The efficiency and scalability of our framework allow the construction of a comprehensive magnetic-structure database. In the present work, we focus on compounds with propagation vector $\mathbf{q}=0$ and fewer than 20 atoms per primitive cell as a starting point. Extensions to larger unit cells and to magnetic structures with $\mathbf{q}\neq 0$ are straightforward within our framework and are currently underway. Moreover, applying our strategy to two-dimensional systems is particularly timely and significant, given the rapidly growing interest in low-dimensional magnetism. Beyond the topological classification already established in this database, we are actively exploring additional magnetic-related physical properties, including multiferroic responses, anomalous Hall conductivity, transport coefficients, etc.